OpenStudy (anonymous):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 9 inches. The height of the cone is 18 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (anonymous):

volume of a cone= \[\pi r^2 \frac{ h }{ 3 }\] volume of a cylinder = \[\pi r^2 h\] so using those equations we can see cylinder volume = 452.39 and the cone volume = 301.59. the relationship of these two objects is a ratio of 2(cone):3(cylinder).

OpenStudy (jack1):

well explained @aay3b